OpenStudy (anonymous):

Solve and show work! (1/x) = (6/3x) + 1

OpenStudy (anonymous):

I'm getting tired of doing peoples homeworks... 1/x = 2/x +1 -1/x = 1 x=-1

OpenStudy (anonymous):

(1/x) = (6/3x) + 1 3/3x=6/3x + 3x/3x 3=6+3x -3=3x -3/3=x -1=x

OpenStudy (anonymous):

denominator comum: 3x ;)

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

why doing it easily when it can be done the hard way :D